Find the indicated points.
If (-1,3) is the endpoint of a line segment, and (-5,1) is its midpoint, find the other endpoint.

Answer by mananth(16946)

mid point (x,y)
x= (x1+x2)/2
y= (y1+y2)/2
mid point (-5,1)
endpoint (-1,3)
-5=(x1+(-1))/2
-5*2=x1-1
x1=-9
1=(y1+3)/2
2=y1+3
y1=-1
endpoint = (-9,-1)